# Recovering contact forms from boundary data

### v5

@@ Abstract.
Let $X$ be a compact connected smooth manifold with boundary. The paper deals with contact $1$-forms $\beta$ on $X$, whose Reeb vector fields $v_{\beta}$ admit Lyapunov functions $f$.
@@ 1. Introduction
The earlier version of this paper [K6] contains a wrong claim (Lemma 6.1). In the present version, we have deleted the old Section 6, Proposition 7.2, and Corollary 7.2, which depended on that erroneous lemma. Also the present Conjecture 4.2 was stated in the earlier versions as Theorem 4.1.
@@ Introduction.
    The main results of this section is Corollary 5.2 dealing with special $1$-parameter families of contact forms in the spirit of Gray Stability Theorem [Gray]. In this section, we analyze the restrictions imposed on contact fields and their flows by the existence of flow-invariant boundary $\partial X$ (see Corollary 5.1).
@@
    This section contains the “non-squeezing” Theorem 6.2. It imposes constraints on contact embeddings $\Psi$ of $(2n+1)$-compact manifolds with boundary into a given contact $(2n+1)$-manifold, which carries a traversing Reeb vector field. It is an analogue of Gromov’s non-squeezing Theorem [Gr2] in Symplectic Geometry. The new phenomenon is that the $(2n-1)$-volume of the equatorial locus in the target restricts the $(2n-1)$-volume of the equatorial locus in the source, provided that we restrict a priory the number of times the Reeb trajectories in the target can hit the boundary of the $\Psi$-image.
@@
    Shadows of Legandrian submanifolds of $(X,\beta)$ on the screen $\partial_{1}^{+}X(v_{\beta})$ and on the trajectory space $\mathcal{T}(v_{\beta})$.
    In this section, we consider Legandrian submanifolds $L$ of a given contact compact manifold $X$ whose Reeb vector field $v_{\beta}$ is traversing. Proposition 7.1 describes the structures on the boundary $\partial X$, sufficient for reconstructing $L$. These structures include the $(-v_{\beta})$-flow generated Lagrangian “shadow” $L^{\dagger}\subset\partial X$ of $L$. For sub-Legandrian links, in Corollary 7.2, we give a boundary-localized criterion when the $v_{\beta}$-flow through $L$ interacts with a $v_{\beta}$-concave portion of the boundary $\partial X$.
@@ 2. Traversing, boundary generic, & traversally generic vector fields
In his groundbreaking work [Mo], M. Morse made an important observation: for a generic vector field $v$, the tangent locus $\partial_{2}X(v)$ inherits a similar structure in connection to $\partial_{1}^{+}X(v)$, as $\partial_{1}X$ has in connection to $X$. That is, $v$ gives rise to a partition $\partial_{2}^{+}X(v)\cup\partial_{2}^{-}X(v)$ of $\partial_{2}X(v)$ into two sets: the locus $\partial_{2}^{+}X(v)$, where the vector field $v$ is directed inward of $\partial_{1}^{+}X(v)$ or is tangent to $\partial_{2}X(v)$, and $\partial_{2}^{-}X(v)$, where it is directed outward of $\partial_{1}^{+}X(v)$ or is tangent to $\partial_{2}X(v)$. Again, we assume that the restriction $v|_{\partial_{2}X(v)}$, viewed as a section of the quotient line bundle $T(\partial X)/T(\partial_{2}X(v))$ over $\partial_{2}X(v)$, is transversal to its zero section.
@@
We will often use the notation “$\partial_{j}^{\pm}X$” instead of “$\partial_{j}^{\pm}X(v)$” when the vector field $v$ is fixed or its choice is obvious.
These considerations motivate a more formal definition.
@@
For the rest of the paper, we assume that the field $v$ on $X$ extends to a vector field $\hat{v}$ on some open manifold $\hat{X}$ which properly contains $X$. We treat the extension $(\hat{X},\hat{v})$ as a germ $Op(X,v)$ that contains $(X,v)$. One may think of $\hat{X}$ as being obtained from $X$ by attaching an external collar to $X$ along $\partial_{1}X$. In fact, the treatment of $(X,v)$ will not depend on the germ of extension $(\hat{X},\hat{v})$, but many constructions are simplified by introducing an extension.
The trajectories $\gamma$ of a boundary generic vector field $v$ on $X$ interact with the boundary $\partial_{1}X$ so that each point $a\in\gamma\cap\partial_{1}X$ acquires a multiplicity $m(a)\in\mathbb{N}$, the order of tangency of $\gamma$ to $\partial_{1}X$ at $a$. We associate a divisor $D_{\gamma}=\sum_{a\,\in\,\gamma\cap\partial_{1}X}m(a)\cdot a$ with each $v$-trajectory $\gamma$. In fact, for any boundary generic $v$, $m(a)\leq\dim(X)$ and the support of $D_{\gamma}$ is finite ([K2], Lemma 3.1).
@@
We consider also an important subclass of traversing and boundary generic fields $v$, which we call traversally generic (see Definition 2.3 below and Definition 3.2 from [K2]). Such vector fields admit $v$-flow-adjusted local coordinate systems, in which the boundary is given by a quite special polynomial equations, and all the trajectories are parallel to the preferred coordinate axis (see [K2], Lemma 3.4). More importantly, for such vector fields, the smooth topological type of the flow in the vicinity of each trajectory of a combinatorial type $\omega$ is determined by $\omega$!
@@ 3. Holography on manifolds with boundary and causality maps
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold with boundary and $v$ a smooth traversing vector field. Recall that $v$ admits a Lyapunov function $f:X\to\mathbb{R}$ so that $df(v)>0$. We assume that $v$ is boundary generic as in Definition 2.2.
@@
that takes any point $x\in\partial_{1}^{+}X(v)$ to the unique consecutive point $y\in\gamma_{x}\cap\partial_{1}^{-}X(v)$, $y\neq x$, which can be reached from $x$ in the $v$-direction. If no such $y$ is available, we put $C_{v}(x)=x$. We stress that typically $C_{v}$ is a discontinuous map.
@@
(3.2) $(\partial_{1}X,\;\succ_{v},)\;$${{\longrightarrow}}$ $\;(X,\;\mathcal{F}(v)),$
(3.3) $(\partial_{1}X,\;\succ_{v},\;f^{\partial})\;$${{\longrightarrow}}$ $\;(X,\;\mathcal{F}(v),\;f),$
where $\succ_{v}$ denotes the partial order on boundary, defined by the causality map $C_{v}$, and the symbol “${{\longrightarrow}}$” indicates the unknown ingredients of the diagram.
The second kind of problem is:
@@
[…]
@@ 4. Contact Holography: recovering contact geometry from the boundary data
Given a vector $v$ and an exterior $k$-form $\alpha$, in what follows, we denote by $v\,\rfloor\,\alpha$ the $(k-1)$-form $\alpha(v\wedge\sim)$.
@@
[…]
@@
Using that each $v$-trajectory $\gamma$ extends to a closed segment of the unique $\hat{v}$-trajectory $\hat{\gamma}$, we construct a tubular cover $\{U_{\alpha}\subset\hat{X}\}_{\alpha}$ of $X$ such that there are diffeomorphisms $\phi_{\alpha}:D^{d-1}\times D^{1}\to U_{\alpha}$ that map each segment $a\times D^{1}$, $a\in D^{d-1}$, to a segment of a $\hat{v}$-trajectory. Moreover, $\phi_{\alpha}(D^{d-1}\times\partial D^{1})\subset\hat{X}\setminus X$. Using compactness of $X$, we choose a finite subcover of the cover $\{U_{\alpha}\}_{\alpha}$. Abusing notations, we denote it $\{U_{\alpha}\}_{\alpha}$.
@@
where $a\in D^{d-1}$. Note that $dg_{\alpha}|_{\gamma}=\hat{\beta}|_{\gamma}$, where $\gamma=\{\phi_{\alpha}(a\times t)\}_{t}$. Since $\hat{\beta}(\hat{v})=1$, we get $dg_{\alpha}(\hat{v})=1$. The finite cover $\{U_{\alpha}\}_{\alpha}$ gives rise to a finite cover $\{V_{\alpha}\}_{\alpha}$ of the compact trajectory space $\mathcal{T}(v)$. A function $\psi:\mathcal{T}(v)\to\mathbb{R}$ is defined to be smooth, if its pull-back $\Gamma^{\ast}(\psi):X\to\mathbb{R}$ is smooth. We use a smooth partition of unity $\{\psi_{\alpha}:V_{\alpha}\to[0,1]\}_{\alpha}$ to introduce a smooth function $f=_{\mathsf{def}}\sum_{\alpha}\psi_{\alpha}\cdot g_{\alpha}.$ By its construction, $\psi_{\alpha}$ has the property $d\psi_{\alpha}(v)=0$. Therefore,
@@
By a similar formula, any closed interval $[a,b]\subset\gamma$, we get
[…]
@@
A coorientable contact structure on a $(2n+1)$-dimensional smooth manifold $X$ is a $2n$-dimensional distribution $\xi$ of cooriented hyperplanes in the tangent bundle $TX$, generated by the kernels of a differential $1$-form $\beta$ on $X$ such that $\beta\wedge(d\beta)^{n}>0$ everywhere. Equivalently, the restriction $(d\beta)^{n}|_{\xi}>0$ everywhere.
An coorientable almost contact structure $\xi$ on a $(2n+1)$-dimensional smooth manifold $X$ is a $2n$-dimensional distribution $\eta$ of cooriented hyperplanes in the tangent bundle $TX$, generated by the kernels of a differential $1$-form $\beta$ on $X$, a nonvanishing vector field $v$ such that $\beta(v)=1$ (so, $v$ is transversal to the distribution $\eta$), and a nondegenerate skew-symmetric $2$-form $\omega$ on $\eta$ (equivalently, a complex structure $J:\eta\to\eta$ such that $J^{2}=-Id$).
@@
A vector field $v_{\beta}$ on $X$ is called the Reeb field for a contact $1$-form $\beta$, if $\beta(v_{\beta})=1$ and $v_{\beta}\in\mathsf{ker}(d\beta)$. $\diamondsuit$
The property $\beta\wedge(d\beta)^{n}\neq 0$ implies that the kernel $\mathsf{ker}(d\beta)$ is $1$-dimensional. Thus, any contact form $\beta$ generates a unique Reeb vector field $v_{\beta}$.
@@
For a contact structure $\xi$, generated by a contact $1$-form $\beta$, there is a $1$-to-$1$ correspondence between the space $C^{\infty}(X)$ of smooth functions and the space $\mathcal{R}eeb(X,\xi)$ of Reeb vector fields amenable to $\xi$. The correspondence is given by the formula $f\longrightarrow v_{e^{f}\cdot\beta}$.
@@
Let us fix a contact form $\beta$ such that $\ker(\beta)=\xi$. Then any other contact form, producing $\xi$ can be written as $e^{f}\beta$ for some smooth function $f$. By the definition of Reeb fields, we get $e^{f}\beta(v_{e^{f}\beta})=1=e^{g}\beta(v_{e^{g}\beta})$. If $v_{e^{f}\beta}=v_{e^{g}\beta}$, this is equivalent to the equation $e^{f}=e^{g}$, or to $f=g$. Therefore, the correspondence $f\longrightarrow v_{e^{f}\cdot\beta}$ is a continuous bijection. ∎
@@
Then $(d\beta)^{n}|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}>0$, while $(d\beta)^{n}|_{\mathsf{int}(\partial_{1}^{-}X(v_{\beta}))}<0$. Thus, $d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}$ is a symplectic $2$-form.
@@
Since $(d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))})^{n}>0$, by Stokes’ Theorem,
@@
Using that each $v$-trajectory $\gamma$ extends to a closed segment of the unique $\hat{v}$-trajectory $\hat{\gamma}$, we construct a finite tubular cover $\{U_{\alpha}\subset\hat{X}\}_{\alpha}$ of $X$ as in the proof of Lemma 4.1. The finite cover $\{U_{\alpha}\}_{\alpha}$ gives rise to a finite cover $\{V_{\alpha}\}_{\alpha}$ of the compact trajectory space $\mathcal{T}(v)$.
@@
Using the diffeomorphism $\phi_{\alpha}:D^{2n}\times D^{1}\to U_{\alpha}$ as in the proof of Lemma 4.1, we transfer $\rho$ to the $1$-form $\rho_{\alpha}=_{\mathsf{def}}\phi_{\alpha}^{\dagger}(\rho)$ and the $1$-form $dx_{2n+1}$ to $\phi_{\alpha}^{\dagger}(dx_{2n+1})=df$ on $U_{\alpha}$. Put $\beta_{\alpha}=_{\mathsf{def}}df+\rho_{\alpha}$.
@@
We use a smooth partition of unity $\{\psi_{\alpha}:V_{\alpha}\to[0,1]\}_{\alpha}$ by smooth functions on $\mathcal{T}(v)$ (i.e., by $v$-invariant smooth functions $\psi_{\alpha}$ on $X$) to introduce a smooth $1$-form $\beta=_{\mathsf{def}}\sum_{\alpha}\psi_{\alpha}\beta_{\alpha}$. By its construction, $\psi_{\alpha}$ has the property $d\psi_{\alpha}(v\wedge\sim)=0$. Therefore,
@@
Thus, $\beta$ is a non-vanishing $1$-form. At the same time, since $d(df)=0$, $\psi_{\alpha}(v)=0$, $v\,\rfloor\,d\psi_{\alpha}=0$, and $\rho_{\alpha}(v)=0$, $v\,\rfloor\,d\rho_{\alpha}=0$, we get:
@@
$=v\,\rfloor\,\big{(}\sum_{\alpha}d\psi_{\alpha}\wedge\rho_{\alpha}\big{)}+v\,\rfloor\,\big{(}\sum_{\alpha}\psi_{\alpha}\,d\rho_{\alpha}\big{)}=0.$
If we would know that $\beta\wedge(d\beta)^{n}>0$ in $X$, we would conclude that $\beta$ is a contact form and $v$ is its Reeb vector field. However, the property $\beta\wedge(d\beta)^{n}>0$ does not hold automatically. Proposition 4.1 spells out the hypotheses that insure the desired property.
@@
If $\partial_{2}^{-}X(v)=\emptyset$ (i.e., if $v$ is concave), then
@@
if and only if,
$(d\beta)^{n}\big{|}_{\mathsf{int}(\partial_{1}^{+}X(v))}>0.$
@@
Let $w_{1},w_{2},\ldots w_{2n}$ be a frame in the tangent space $T_{x}\partial_{1}^{+}X(v)$ whose orientation agrees with the orientation of $\partial_{1}X(v)$, induced by the preferred orientation of $X$. Then $v\wedge w_{1}\wedge w_{2}\wedge\ldots\wedge w_{2n}$ is a positive volume. Thus, $(\beta\wedge(d\beta)^{n})(v\wedge w_{1}\wedge w_{2}\wedge\ldots\wedge w_{2n})>0$ on $\mathsf{int}(\partial_{1}^{+}X(v))$. Using that $\beta$ is $v$-invariant, we conclude that $\beta\wedge(d\beta)^{n}$ is a positive volume form along each $v$-trajectory $\gamma$ such that $\gamma\cap\mathsf{int}(\partial_{1}^{+}X(v))\neq\emptyset$. If $\partial_{2}^{-}X(v)=\emptyset$ (this implies that $\partial_{j}X(v))=\emptyset$ for all $j\geq 3$), any trajectory $\gamma$ has this property. Therefore, in such a case, we conclude that $\beta\wedge(d\beta)^{n}>0$ everywhere in $X$. As a result, $\beta$ is a contact form with the Reeb vector field $v_{\beta}=v$ and the bundle $\xi_{\beta}\approx\xi(v)$ admits a symplectic structure, induced by $d\beta$ (equivalently, a complex structure $J:\xi_{\beta}\to\xi_{\beta}$, consistent with $d\beta$).

### v6

@@ Abstract.
Let $X$ be a compact smooth manifold with boundary. The paper deals with contact $1$-forms $\beta$ on $X$, whose Reeb vector fields $v_{\beta}$ admit Lyapunov functions $f$.
@@ 1. Introduction
Erratum. The earlier version [K6] of this paper contains a wrong claim, Corollary 4.7. In the present version, it is replaced by Lemma 4.6. Also, the old Corollary 5.2, based on the erroneous Corollary 4.7, is removed. $\diamondsuit$
@@ Introduction.
    In this section, we analyze the restrictions imposed on contact fields and their flows by the existence of flow-invariant boundary $\partial X$ (see Theorem 5.1).
@@
    This section contains “contact non-squeezing” Theorem 6.2. It imposes constraints on contact embeddings $\Psi$ of $(2n+1)$-compact manifolds $X$ with boundary into a given contact $(2n+1)$-manifold $Y$, which carries a traversing Reeb vector field. It is an analogue of Gromov’s non-squeezing Theorem [Gr2] in Symplectic Geometry. The new phenomenon is that the (2n1)(2n-1)-volume of the equatorial locus in the target restricts the (2n1)(2n-1)-volume of the equatorial locus in the source, provided that we restrict a priory the number of times the Reeb trajectories in the target can hit the boundary of the $\Psi$-image.
@@
    Shadows of Legendrian submanifolds of $(X,\beta)$ on the screen $\partial_{1}^{+}X(v_{\beta})$ and on the trajectory space $\mathcal{T}(v_{\beta})$.
    In this section, we consider Legendrian submanifolds $L$ of a given contact compact manifold $X$ whose Reeb vector field $v_{\beta}$ is traversing. Proposition 7.1 describes the structures on the boundary $\partial X$, sufficient for reconstructing $L$. These structures include the $(-v_{\beta})$-flow generated Lagrangian “shadow” $L^{\dagger}\subset\partial X$ of $L$.
    In Corollary 7.2, for sub-Legendrian links, we give a boundary-localized criterion when the $v_{\beta}$-flow through $L$ interacts with a $v_{\beta}$-concave portion of the boundary $\partial X$.
@@ 2. Traversing, boundary generic, & traversally generic vector fields
In his groundbreaking work [M], M. Morse made an important observation: for a generic vector field $v$, the tangent locus $\partial_{2}X(v)$ inherits a similar structure in connection to $\partial_{1}^{+}X(v)$, as $\partial_{1}X$ has in connection to $X$. That is, $v$ gives rise to a partition $\partial_{2}^{+}X(v)\cup\partial_{2}^{-}X(v)$ of $\partial_{2}X(v)$ into two sets: the locus $\partial_{2}^{+}X(v)$, where the vector field $v$ is directed inward of $\partial_{1}^{+}X(v)$ or is tangent to $\partial_{2}X(v)$, and $\partial_{2}^{-}X(v)$, where it is directed outward of $\partial_{1}^{+}X(v)$ or is tangent to $\partial_{2}X(v)$. Again, we assume that the restriction $v|_{\partial_{2}X(v)}$, viewed as a section of the quotient line bundle $T(\partial X)/T(\partial_{2}X(v))$ over $\partial_{2}X(v)$, is transversal to its zero section.
@@
We will often use the notation $\partial_{j}^{\pm}X$” instead of $\partial_{j}^{\pm}X(v)$” when the vector field $v$ is fixed or its choice is obvious.
These considerations motivate a more formal definition ([K1], [K3]).
@@
[…]
The trajectories $\gamma$ of a boundary generic vector field $v$ on $X$ interact with the boundary $\partial_{1}X$ so that each point $a\in\gamma\cap\partial_{1}X$ acquires a multiplicity $m(a)\in\mathbb{N}$, the order of tangency of $\gamma$ to $\partial_{1}X$ at $a$. We associate a divisor $D_{\gamma}=\sum_{a\,\in\,\gamma\cap\partial_{1}X}m(a)\cdot a$ with each $v$-trajectory $\gamma$. In fact, for any boundary generic $v$, the multiplicity $m(a)\leq\dim(X)$ and the support of $D_{\gamma}$ is finite ([K2], Lemma 3.1).
@@
We consider also an important subclass of traversing and boundary generic fields $v$, which we call traversally generic (see Definition 2.3 below and Definition 3.2 from [K2]). Such vector fields admit $v$-flow-adjusted semi-local coordinate systems, in which the boundary is given by a quite special polynomial equations, and all the trajectories are parallel to the preferred coordinate axis (see [K2], Lemma 3.4). More importantly, for such vector fields, the smooth topological type of the flow in the vicinity of each trajectory of a combinatorial type $\omega$ is determined by $\omega$ (see [K1])!
@@ 3. Holography on manifolds with boundary and causality maps
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold with boundary and $v$ a smooth traversing vector field. Recall that $v$ admits a Lyapunov function $f:X\to\mathbb{R}$ so that $df(v)>0$. We assume that $v$ is boundary generic in the sense of Definition 2.2.
@@
(see [K2] or [K4]) that takes any point $x\in\partial_{1}^{+}X(v)$ to the unique consecutive point $y\in\gamma_{x}\cap\partial_{1}^{-}X(v)$, $y\neq x$, which can be reached from $x$ in the $v$-direction. If no such $y$ is available, we put $C_{v}(x)=x$. We stress that typically $C_{v}$ is a discontinuous map.
@@
$\bullet\text{\sf Holographic Reconstruction Problem}$
(3.2) $(\partial_{1}X,\;\succ_{v},)\;$ $\stackrel{{\scriptstyle\mathbf{??}}}$$\longrightarrow$}} $\;(X,\;\mathcal{F}(v)),$
(3.3) $(\partial_{1}X,\;\succ_{v},\;f^{\partial})\;$ $\stackrel{{\scriptstyle\mathbf{??}}}$$\longrightarrow$}} $\;(X,\;\mathcal{F}(v),\;f),$
@@
The second kind of problem is stated as follows:
@@
$\bullet\text{\sf Holographic Extension Problem \quad}$
@@ 4. Contact Holography: recovering contact geometry from the boundary data
Given a vector $v$ and an exterior $k$-form $\alpha$, in what follows, we denote by $v\,\rfloor\,\alpha$ the (k1)(k-1)-form $\alpha(v\wedge\sim)$.
Let $\alpha$ be a differential $k$-form on a smooth manifold $X$ and let $Z\subset X$ be a smooth submanifold. We use somewhat unconventional notations: we write $\alpha|_{Z}$” for the restriction of the section $\alpha:X\to\bigwedge^{k}T^{\ast}X$ to the section $\alpha|:Z\to\bigwedge^{k}T^{\ast}Z$, and we write $\alpha||_{Z}$” for the restriction of $\alpha$ to the section of the bundle $\bigwedge^{k}T^{\ast}X|_{Z}\to Z$.
@@
Let $v$ be a smooth traversing vector field and $\beta$ a $1$-form on a compact manifold $X$ such that $\beta(v)=1$. Then $v$ admits a Lyapunov function $f:X\to\mathbb{R}$ such that $df(v)=1$. Moreover, for any $v$-trajectory $\gamma$ and any closed interval $a,b]\subset\gamma$, we have $\int_{[a,b]}df=\int_{[a,b]}\beta$.
@@
Using that each $v$-trajectory $\gamma$ extends to a closed segment of the unique $\hat{v}-$trajectory $\hat{\gamma}$, we construct a tubular cover $\{U_{\alpha}\subset\hat{X}\}_{\alpha}$ of $X$ such that there are diffeomorphisms $\phi_{\alpha}:D^{d-1}\times D^{1}\to U_{\alpha}$ that map each segment $a\times D^{1}$, $a\in D^{d-1}$, to a segment of a $\hat{v}-$trajectory. Moreover, $\phi_{\alpha}(D^{d-1}\times\partial D^{1})\subset\hat{X}\setminus X$. Using compactness of $X$, we choose a finite subcover of the cover $\{U_{\alpha}\}_{\alpha}$. Abusing notations, we still denote it $\{U_{\alpha}\}_{\alpha}$.
@@
where $a\in D^{d-1}$. Note that $dg_{\alpha}|_{\gamma}=\hat{\beta}|_{\gamma}$, where $\gamma=\{\phi_{\alpha}(a\times t)\}_{t}$. Since $\hat{\beta}(\hat{v})=1$, we get $dg_{\alpha}(\hat{v})=1$. The finite cover $\{U_{\alpha}\}_{\alpha}$ gives rise to a finite cover $\{V_{\alpha}\}_{\alpha}$ of the compact trajectory space $\mathcal{T}(v)$.
A function $\psi:\mathcal{T}(v)\to\mathbb{R}$ is defined to be smooth, if its pull-back $\Gamma^{\ast}(\psi):X\to\mathbb{R}$ is smooth. We use a smooth partition of unity $\{\psi_{\alpha}:V_{\alpha}\to[0,1]\}_{\alpha}$ to introduce a smooth function $f=_{\mathsf{def}}\sum_{\alpha}\psi_{\alpha}\cdot g_{\alpha}.$ By its construction, $\psi_{\alpha}$ has the property $d\psi_{\alpha}(v)=0$. Therefore,
@@
By a similar formula, any closed interval $a,b]\subset\gamma$, we get
$\int_{[a,b]}df=\sum_{\alpha}\Big(\int_{[a,b]}\psi_{\alpha}\cdot dg_{\alpha}+\int_{[a,b]}g_{\alpha}\cdot d\psi_{\alpha}\Big)=\sum_{\alpha}\Big(\psi_{\alpha}\int_{[a,b]}dg_{\alpha}\Big)$
@@
A coorientable contact structure on a $(2n+1)$-dimensional smooth manifold $X$ is a $2n$-dimensional distribution $\xi$ of cooriented hyperplanes in the tangent bundle $TX$, generated by the kernels of a differential $1$-form $\beta$ on $X$ such that $\beta\wedge(d\beta)^{n}>0$ everywhere (equivalently, the restriction $(d\beta)^{n}|_{\xi}>0$ everywhere).
An coorientable almost contact structure $\xi$ on a $(2n+1)$-dimensional smooth manifold $X$ is a $2n$-dimensional distribution $\eta$ of cooriented hyperplanes in the tangent bundle $TX$, generated by the kernels of a differential $1$-form $\beta$ on $X$, a nonvanishing vector field $v$ such that $\beta(v)=1$ (so, $v$ is transversal to the distribution $\eta$), and a nondegenerate skew-symmetric $2$-form $\omega$ on $\eta$ (equivalently, a complex structure $J:\eta\to\eta$ such that $J^{2}=-\mathsf{Id}$).
@@
A vector field $v_{\beta}$ on $X$ is called the Reeb field for a contact $1$-form $\beta$, if $\beta(v_{\beta})=1$ and $v_{\beta}\in\ker(d\beta)$. $\diamondsuit$
The property $\beta\wedge(d\beta)^{n}\neq 0$ implies that the kernel $\ker(d\beta)$ is $1$-dimensional. Thus, any contact form $\beta$ generates a unique Reeb vector field $v_{\beta}$.
@@
For a contact structure $\xi_{\beta}$, generated by a contact $1$-form $\beta$, there is a 11-to-11 correspondence between the space $C^{\infty}(X)$ of smooth functions $f$ and the space $\mathcal{R}eeb(X,\xi_{\beta})$ of Reeb vector fields amenable to $\xi_{\beta}$. The correspondence is given by the formula $f\longrightarrow v_{e^{f}\cdot\beta}$.
@@
Let us fix a contact form $\beta$ such that $\ker(\beta)=\xi_{\beta}$. Then any other contact form, producing $\xi_{\beta}$ can be written as $e^{f}\beta$ for some smooth function $f$. By the definition of Reeb fields, we get $e^{f}\beta(v_{e^{f}\beta})=1=e^{g}\beta(v_{e^{g}\beta})$. If $v_{e^{f}\beta}=v_{e^{g}\beta}$, this is equivalent to the equation $e^{f}=e^{g}$, or to the equation $f=g$. Therefore, the correspondence $f\longrightarrow v_{e^{f}\cdot\beta}$ is a continuous bijection. ∎
@@
Then $(d\beta)^{n}|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}>0$, while $(d\beta)^{n}|_{\mathsf{int}(\partial_{1}^{-}X(v_{\beta}))}<0$. Thus, the restriction $d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}$ is a symplectic $2$-form.
@@
Since, by Lemma 4.3, $(d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))})^{n}>0$, by Stokes’ Theorem,
@@
Using that each $v$-trajectory $\gamma$ extends to a closed segment of the unique $\hat{v}-$trajectory $\hat{\gamma}\subset\hat{X}$, we construct a finite tubular cover $\{U_{\alpha}\subset\hat{X}\}_{\alpha}$ of $X$ as in the proof of Lemma 4.1. The finite cover $\{U_{\alpha}\}_{\alpha}$ gives rise to a finite cover $\{V_{\alpha}\}_{\alpha}$ of the compact trajectory space $\mathcal{T}(v)$.
@@
Using the inverse $\phi_{\alpha}^{\dagger}$ of the diffeomorphism $\phi_{\alpha}:D^{2n}\times D^{1}\to U_{\alpha}$ as in the proof of Lemma 4.1, we transfer the $1$-form $\rho$ to the $1$-form $\rho_{\alpha}=_{\mathsf{def}}\phi_{\alpha}^{\dagger}(\rho)$ and the $1$-form $dx_{2n+1}$ to $\phi_{\alpha}^{\dagger}(dx_{2n+1})=df$ on $U_{\alpha}$. Put $\beta_{\alpha}=_{\mathsf{def}}df+\rho_{\alpha}$.
@@
We use a smooth partition of unity $\{\psi_{\alpha}:V_{\alpha}\to[0,1]\}_{\alpha}$ by smooth functions on $\mathcal{T}(v)$ (i.e., by $v$-invariant smooth functions $\psi_{\alpha}$ on $X$) to introduce a smooth $1$-form $\beta=_{\mathsf{def}}\sum_{\alpha}\psi_{\alpha}\beta_{\alpha}$. By its construction, each $\psi_{\alpha}$ has the property $d\psi_{\alpha}(v\wedge\sim)=0$. Therefore,
@@
Thus, $\beta$ is a non-vanishing $1$-form. At the same time, using that $d(df)=0$, $\psi_{\alpha}(v)=0$, $v\,\rfloor\,d\psi_{\alpha}=0$, and $\rho_{\alpha}(v)=0$, and $v\,\rfloor\,d\rho_{\alpha}=0$, we get:
@@
$v\,\rfloor\,\big(\sum_{\alpha}d\psi_{\alpha}\wedge\rho_{\alpha}\big)+v\,\rfloor\,\big(\sum_{\alpha}\psi_{\alpha}\,d\rho_{\alpha}\big)=0.$
If we would know that $\beta\wedge(d\beta)^{n}>0$ in $X$, we would conclude that $\beta$ is a contact form and $v$ is its Reeb vector field. However, the property $\beta\wedge(d\beta)^{n}>0$ does not hold automatically. Proposition 4.1 spells out some hypotheses that insure the desired property.
@@
Assuming that $\partial_{2}^{-}X(v)=\emptyset$ (i.e., when $v$ is concave), the three properties
@@
are valid, if and only if, $(d\beta)^{n}\big|_{\mathsf{int}(\partial_{1}^{+}X(v))}>0.$ [^1] Thus, if the bundle $\xi(v)$ does not admit a complex structure, then the restriction of the $2n$-form $(d\beta)^{n}$ to $\mathsf{int}(\partial_{1}^{+}X(v))$ must switch sines.
@@
Let $w_{1},w_{2},\ldots w_{2n}$ be a frame in the tangent space $T_{x}\partial_{1}^{+}X(v)$ whose orientation agrees with the orientation of $\partial_{1}^{+}X(v)$, induced by the preferred orientation of $X$. Then $v\wedge w_{1}\wedge w_{2}\wedge\ldots\wedge w_{2n}$ is a positive volume. Thus, $(\beta\wedge(d\beta)^{n})(v\wedge w_{1}\wedge w_{2}\wedge\ldots\wedge w_{2n})>0$ along $\mathsf{int}(\partial_{1}^{+}X(v))$. Using that $\beta$ is $v$-invariant, we conclude that $\beta\wedge(d\beta)^{n}$ is a positive volume form along each $v$-trajectory $\gamma$ such that $\gamma\cap\mathsf{int}(\partial_{1}^{+}X(v))\neq\emptyset$. Thus, in $X\setminus\partial_{2}^{-}X(v)$, the form $\beta\wedge(d\beta)^{n}$ is a positive. If $\partial_{2}^{-}X(v)=\emptyset$ (i.e., the boundary is concave with respect to $v$), we conclude that $\beta\wedge(d\beta)^{n}>0$ everywhere in $X$. As a result, $\beta$ is a contact form with the Reeb vector field $v_{\beta}=v$ and the bundle $\xi_{\beta}\approx\xi(v)$ admits a symplectic structure, induced by $d\beta|_{\xi(v)}$ (equivalently, a complex structure $J:\xi_{\beta}\to\xi_{\beta}$, consistent with $d\beta$).